Circular No. 8209

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 U 1 AND S/2003 U 2 M. R. Showalter, Stanford University; and J. J. Lissauer, Ames Re-

search Center, NASA, report the discovery of two previously unknown inner satellites of Uranus in recent data from the High Resolution Channel of the Hubble Space Telescope (HST) Advanced Camera for Surveys. A set of twenty-four 250-s clear-filter (bandpass 300–900 nm) exposures was obtained during three consecutive HST orbits between Aug. 25.8042 and 25.9659 UT. Two additional 120-s images were obtained during July 10.7273–10.7436, and two 200-s images were obtained during Sept. 11.4765– 11.5059. S/2003 U 1 is visible on sixteen images from Aug. 25 and on all images from July 10 and Sept. 11; its mean motion is 390^{o}.06 ± 0^{o}.02 day^{−1}, corresponding to a semimajor axis of 97734 ± 4 km, and it appeared at maximum elongation off the northern ring ansa on Aug. 25.97. S/2003 U 2 is visible in six images obtained during Aug. 25.8042–25.8297; it crossed the southern ansa on Aug. 25.823 at radial distance 74800 ± 200 km, and its motion is consistent with a circular orbit at this distance. S/2003 U 1 is ≈ 5 percent as bright as Uranus XIV (Belinda), with a signal-to-ratio (S/N) of 5–10 in most individual images, while S/2003 U 2 is ≈ 3 percent as bright as Uranus XIV, with S/N = 4 in individual images.

COMET P/2003 S2 (NEAT) Further to IAUC 8208, M. Hicks reports the NEAT discovery of another

comet on images obtained at Haleakala (discovery observation below), noting a coma diameter of 5′′ and a 15′′ tail in p.a. 255^{o}. Following posting on the NEO Confirmation Page, numerous observers have reported cometary activity from CCD images obtained on Sept. 24.9–25.0 UT, including R. Apitzsch (Wildberg, Germany; 2′′ coma), P. Birtwhistle (not ‘Birthwhistle’, as given on IAUC 8208; 8′′ coma with a broad tail at least 20′′ long in p.a. 250^{o} and a thin tail 20′′ long in p.a. 330^{o}), F. Hormuth (Heppenheim, Germany; 40′′ tail in p.a. 260^{o}), and Š. Gajdoš and A. Galád (Modra, Slovakia; coma diameter ∼ 15′′ and tail extending in p.a. 255^{o}–260^{o}). The available astrometry (including prediscovery LINEAR observations from Sept. 19.4), preliminary orbital elements (T = 2003 July 10.0 TT, ω = 259^{o}.6, Ω = 91^{o}.5, i = 7^{o}.3, equinox 2000.0, q = 2.294 UT, e = 0.454, P = 8.6 yr), and an ephemeris appear on MPEC 2003-S65.

2003 UT  α_{2000}  δ_{2000}  Mag.
Sept.24.60807  2^{h}00^{m}41.^{s}55  −0^{o}03′43′′.4  18.0
2003 September 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green